A 39 kW delta-connected motor with a lagging power factor of 0,85 is supplied by a three­phase star-connected alternator having a line output of 380 V at 50 Hz. The motor has an efficiency of 78 %. Calculate the phase currents for the load and the alternator. Determine the new line current of the motor, if the supply frequency is changed to 60 Hz


1
Expert's answer
2021-10-27T07:01:42-0400

"\mu*P=\sqrt3VIcos\theta"

"0.78*39*10^3=\sqrt3*380*0.85*I"

"I=\frac{0.78*39*10^3}{\sqrt3*380*0.85}"

=54.3A

For star load, "I_l=I_{ph}"

Thus alternate phase current=54.3A

For delta load, "I_{ph}=\frac{I_l}{\sqrt3}"

Hence motor phase current="\frac{54.3}{\sqrt3}"

=31.35A

When frequency of supply is increased, the flux density and torque decreases.

As the flux density decrease while speed increase they both compensate each other and the power and current remains same.

"I_{ph}=31.35A"